# Any-Time Regret-Guaranteed Algorithm for Control of Linear Quadratic Systems

### v2

@@ Abstract
In this paper, we propose a computationally efficient algorithm that achieves anytime regret of order $\mathcal{O}(\sqrt{t})$, with explicit dependence on the system dimensions and on the solution of the Discrete Algebraic Riccati Equation (DARE). Our approach builds on the SDP-based framework of cohen2019learning, using an appropriately tuned regularization and a sufficiently accurate initial estimate to construct confidence ellipsoids for control design. A carefully designed input-perturbation mechanism is incorporated to ensure anytime performance. We develop two variants of the algorithm. The first enforces strong sequential stability, requiring each policy to be stabilizing and successive policies to remain close. This sequential condition helps prevent state explosion at policy update times and enables high-probability guarantees on the state trajectory; however, it results in a conservative design and suboptimal regret scaling with respect to the DARE solution. Motivated by this limitation, we introduce a second class of algorithms that removes the sequential-stability requirement and instead requires only that each generated policy be stabilizing. Closed-loop stability is then preserved through a dwell-time–inspired policy-update rule, adapting ideas from switched-systems control to carefully balance exploration and exploitation. This class of algorithms also addresses key shortcomings of most existing approaches—including certainty-equivalence–based methods—which typically guarantee stability only in the Lyapunov sense and lack explicit uniform high-probability bounds on the state trajectory expressed in system-theoretic terms. In particular, such methods fail to prevent transient state spikes caused by transitions between stabilizing policies, a critical issue in safety-sensitive systems. Our analysis explicitly characterizes the trade-off between state amplification and regret, and shows that partially relaxing the sequential-stability requirement yields optimal regret. Finally, our method eliminates the need for any a priori bound on the norm of the DARE solution—an assumption required by all existing computationally efficient optimism in the face of uncertainty (OFU) based algorithms—and thereby removes the reliance of regret guarantees on such external inputs.
@@ 1 Introduction
where matrices $A_{*}\in\mathbb{R}^{n\times n}$ and $B_{*}\in\mathbb{R}^{n\times m}$ and the process noise $\omega_{t+1}$ usually is assumed to ben drawn i.i.d from a zero-mean gaussian distribution.
@@
With full knowledge of system parameters $A_{*}$, $B_{*}$, and assuming controllability of $(A_{*},B_{*})$, it is known that the optimal sequence of control inputs $u^{*}=\{u^{*}_{t}\}_{t=1}^{T-1}$ minimizing average expected cost
$J_{\mathcal{A}}=$ $\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}[\sum_{t=0}^{T-1}c_{t}]$ (3)
@@
In the absence of knowledge of the parameters $A_{*}$ and $B_{*}$, the core challenge is to design an algorithm that stabilizes the system while minimizing the accumulated quadratic cost, quantified by a finite-time performance measure. In this study, we adopt regret as the performance measure, defined as follows. For an arbitrary algorithm $\mathcal{A}$, which generates the input sequence $u=\{u_{t}\}_{t=0}^{T-1}$, regret is given by
$R_{\mathcal{A}}(T)=\sum_{t=0}^{T-1}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}-J_{*}\right),$ (4)
@@ 1.1 Contributions and informal statements of results
The first algorithm proposed for the control of LQR with a worst-case regret bound of $\mathcal{O}(\sqrt{T})$ is the Optimism in the Face of Uncertainty (OFU)-based algorithm introduced by abbasi2011regret. This algorithm performs system identification by constructing confidence ellipsoids and designs the control policy by playing optimistically with respect to this set. Although the algorithm enjoys strong regret guarantees, its nonconvexity and the resulting computational burden has spurred subsequent research into more efficient formulations. with guaranteed $\mathcal{O}(\sqrt{T})$.
@@
Another property that may be important for making learning-based algorithms transferable across control domains is the provision of system-theoretic guarantees on the state-norm trajectory. Establishing such bounds, however, requires a thorough analysis of the closed-loop stability induced by the proposed algorithms. Generating individually stabilizing policies is not sufficient to ensure closed-loop stability, as instability (or state explosion) may still occur due to policy updates, a well known fact in literature on switched-systems control. Previous works have attempted to address this challenge through various approaches. For instance, certainty-equivalence (CE) methods simchowitz2020naive; jedra2022minimal employ a common Lyapunov function to guarantee Lyapunov stability of the closed-loop system. However, it is important to note that Lyapunov decay alone does not necessarily provide explicit high-probability, uniform control of the state norm in the presence of process noise. Under zero-mean Gaussian disturbances, Lyapunov decay guarantees only that the expected state norm, $\mathbb{E}\left[\|x_{t}\|^{2}\right]$, remains bounded, and deriving a high-probability bound requires additional effort. Although a high-probability state-norm bound is provided by simchowitz2020naive (see their Lemma 5.3 and its proof), this bound is conservative and does not accurately capture the true scale of possible state-norm excursions. Therefore, while Lyapunov stability is sufficient for controlling accumulated regret—since a finite number of such spikes can typically be absorbed in regret bounds—it is insufficient for obtaining high-probability guarantees on the state norm, which are essential for safety-critical systems. Another work that provides a high-probability state-norm bound is abeille2020efficient, but the bound is not fully system-theoretic and depends on a priori bounds, which may be loose. In another work, cohen2019learning introduced the notion of strong sequential stability, which enables an explicit high-probability guarantee on the state norm. This notion imposes a stricter requirement than standard stability: the algorithm must not only output stabilizing policies, but also ensure that any two consecutively generated policies remain close in a suitable sense, preventing the state norm from blowing up due to policy updates. While this approach provides a high-probability bound on the state norm, in this work we show that the strong sequential-stability requirement is conservative from the perspective of regret guarantees, leading to an $\mathcal{O}(|P_{*}|^{8})$ dependence on the operator norm of the DARE solution. This issue motivates us to propose a method that incorporates a mechanism inspired by dwell-time design in switched-systems control. Using this mechanism, we carefully regulate the exploration–exploitation trade-off to improve regret bounds while explicitly restricting state amplification and ensuring uniform control of the closed-loop trajectories. A natural first idea would be to require the algorithm to generate only stabilizing policies and then rely on a dwell-time condition to control both the state norm and the regret. However, sequential stability is not a binary property; rather, it forms a tunable continuum. Our analysis shows that fully enforcing sequential stability leads to suboptimal regret, while discarding it entirely and relying solely on a dwell-time–type design for exploration–exploitation tuning also degrades regret. The optimal performance arises at an intermediate relaxation level, where stability and exploration are balanced. We identify a specific tuning that guarantees an $\mathcal{O}(\|P_{*}\|^{6})$ dependence in the regret bound, while maintaining control of the state norm with a dependence of order $\mathcal{O}(\|P_{*}\|^{2.5})$. Our analysis reveals a clear trade-off between the state-norm bound and the regret bound in terms of their dependence on $\|P_{*}\|$, and it provides the flexibility to tune this trade-off as needed.
@@
Our core contributions are briefly threefold: (i) We introduce new computationally efficient anytime-regret algorithms for LQ control, motivated by the need for methods that remain effective when the implementation horizon is not known a priori. (ii) We emphasize the importance of providing a high-probability, system-theoretic bound on the state norm. We offer a new theoretical understanding of strong sequential stability, highlight its limitations, and demonstrate how relaxing this notion can improve algorithm performance. A key component of this development is a novel mechanism, inspired by dwell-time design in switched-systems control, which enables principled tuning of the exploration–exploitation trade-off while ensuring both favorable regret bounds and closed-loop stability. (iii) We eliminate the need for a priori knowledge of the optimal average cost $J_{*}$, an assumption required by existing computationally efficient OFU-based algorithms.
@@
$(1)$ Any-time regret algorithm with guaranteed system theoretic regret bound We propose an any-time SDP-based LQ cOntrol (ARSLO) built upon the SDP-based algorithm proposed by cohen2019learning, OSLO which is a horizon dependant algorithm. The OSLO algorithm when running for $T$ time steps requires $T$-dependant regularization parameters used for least square estimation and confidence bound construction as well as an initial parameter estimate in $\mathcal{O}(1/T^{1/4})$ neighborhood of true but unknown system dynamic parameters. The latter requires running a warm-up algorithm for $\mathcal{O}(\sqrt{T})$ time steps. To extend to a longer horizon using the doubling trick, such a warm-up phase must be implemented for the extended version in each epoch. Demonstrating this issue, we propose an any-time regret guaranteed algorithm by choosing an appropriate regularization parameter for the least square estimation and applying additive input perturbation $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, where $\|\Gamma_{t}\|_{op}=\mathcal{O}(1/\sqrt{t})$. Therefore, our ARSLO algorithm, provides anytime regret guarantees while also ensuring the strong sequential stability property of the generated policies (see Definition 1). Specifically, ARSLO achieves the regret bound
@@
Leaving aside the anytime regret guarantee of our proposed algorithms, their system-theoretic regret upper bound challenges that of the Lag-LQR approach proposed by abeille2020efficient. Lag-LQR is developed by applying techniques from extended value iteration (EVI) to the Optimistic LQR with relaxed constraints, followed by a Lagrangian relaxation to design a computationally efficient algorithm. This algorithm comes with the following regret upper bound:
@@ 1.2 Related works
In recent years, the machine learning community has shown increasing interest in addressing control design in the context of LQR, under various scenarios of what is known and what is unknown. Notably, agarwal2019online studied the setting with (nonstochastic) adversarial perturbations and established regret guarantees. cohen2018online investigated the control of linear systems with a known model but adversarially chosen quadratic costs, also providing regret guarantees. When the system model matrices are unknown, which is the main focus of this paper, existing efforts can be categorized into model-free and model-based settings. Model-free approaches design controllers without explicit system identification; instead, they minimize the cost function directly, relying solely on available data (see dean2020sample; fazel2018global; abbasi2019model). On the other hand, model-based approaches perform control design through explicit model identification or confidence set construction and typically provide regret guarantees. Below, we briefly review several relevant papers in the literature, presented in chronological order.
The first algorithm proposed for the control of LQR with a worst-case regret bound of $\mathcal{O}(\sqrt{T})$ is the OFU-based algorithm introduced by abbasi2011regret. This algorithm performs system identification by constructing confidence ellipsoids and designs the control policy by playing optimistically with respect to this set. Despite its success, the algorithm’s nonconvexity became a key motivation for subsequent work in the literature to develop fully convex algorithms with guaranteed $\mathcal{O}(\sqrt{T})$ worst-case regret.
@@
A commonly used categorization of regret-guaranteed algorithms for LQ control—also adopted in jedra2022minimal—classifies existing works into (i) self-tuning regulator–based approaches and (ii) OFU–based approaches. Self-tuning regulator algorithms estimate the unknown system matrices $(A_{*},B_{*})$, treat these estimates as the true parameters, solve the corresponding DARE, and design the control input accordingly. To ensure sufficient excitation of the system—and hence enable consistent estimation of $(A_{*},B_{*})$—the control inputs are typically perturbed with carefully tuned noise. OFU-based approaches, on the other hand, construct a confidence set for the unknown system parameters and design the control policy by acting optimistically with respect to this set. Our algorithm, while OFU-based, also employs input perturbations for exploration. As a result, our approach can be viewed as lying at the intersection of these two categories. A more detailed discussion and quantitative comparisons are provided in Appendix A.
@@ 2 Problem Statement and Preliminaries
Notation. For matrix $M$, $\|M\|_{*}=\operatorname{tr}(\sqrt{M^{\top}M})$ is the trace norm, $\|M\|_{F}=\sqrt{\operatorname{tr}(M^{\top}M)}$ is the Frobenius norm and $\|M\|$ is the spectral norm. $\lambda_{\min}(.)$ represents the minimum eigenvalue. Finally, we use $A\bullet B$ to denote the entry-wise dot product between matrices, namely $A\bullet B=\operatorname{Tr}(A^{\top}B)$. We write $f\lesssim g$ to denote that $f(x)\leq C\,g(x)$ for some universal constant $C>0$. The notation $\mathcal{O}$ indicates dependencies only on the time $t$, the confidence parameter $\delta$, $\|P_{*}\|$, and the dimensions $n$ and $m$.
@@ 2.2 Preliminaries
Confidence Ellipsoid Construction To begin, we outline how a confidence ellipsoid around the true but unknown parameters $\Theta_{*}$ is constructed using the available measurements, leaving further details to the supplementary material. Let $X_{t}$ and $Z_{t}$ denote the matrices whose rows are $x_{1}^{\top},\ldots,x_{t}^{\top}$ and $z_{0}^{\top},\ldots,z_{t-1}^{\top}$, respectively. By applying least-square estimation, regularized with $\operatorname*{Tr}((\Theta-\Theta_{0})^{\top}(\Theta-\Theta_{0}))$ for some $\Theta_{0}$ and regularization parameter $\lambda$, the least-square estimate of $\Theta_{*}$ is given by:
@@
where we denote its optimal solution by $P(\mathcal{C}_{t})$. Note that in cohen2018online, the term $\mu_{t}\|P\|_{*}V_{t}^{-1}$ on the right-hand side of their relaxed dual SDP introduces unnecessary looseness in the analysis and it can be replaced by $\mu_{t}\|P\|V_{t}^{-1}$. Our perturbation lemma (Lemma 7 in Appendix E), together with the corresponding lemma in their analysis, shows that such an SDP relaxation is not necessary.

### v3

@@ Abstract
We propose a computationally efficient algorithm that achieves anytime regret of order $\mathcal{O}(\sqrt{t})$, with explicit dependence on the system dimensions and on the solution of the Discrete Algebraic Riccati Equation (DARE). Our approach builds on the SDP-based framework of cohen2019learning, using an appropriately tuned regularization and a sufficiently accurate initial estimate to construct confidence ellipsoids for control design. A carefully designed input-perturbation mechanism is incorporated to ensure anytime performance. We develop two variants of the algorithm. The first enforces a notion of strong sequential stability, requiring each policy to be stabilizing and successive policies to remain close. However, enforcing this notion results in a suboptimal regret scaling. The second class of algorithms removes the sequential-stability requirement and instead requires only that each generated policy be stabilizing. Closed-loop stability is then preserved through a dwell-time–inspired policy-update rule, adapting ideas from switched-systems control to carefully balance exploration and exploitation. This class of algorithms also addresses key shortcomings of most existing approaches—including certainty-equivalence–based methods—which typically guarantee stability only in the Lyapunov sense and lack explicit uniform high-probability bounds on the state trajectory expressed in system-theoretic terms. Our analysis explicitly characterizes the trade-off between state amplification and regret, and shows that partially relaxing the sequential-stability requirement yields optimal regret. Finally, our method eliminates the need for any a priori bound on the norm of the DARE solution—an assumption required by all existing computationally efficient optimism in the face of uncertainty (OFU) based algorithms—and thereby removes the reliance of regret guarantees on such external inputs.
@@ 1 Introduction
where matrices $A_{*}\in\mathbb{R}^{n\times n}$ and $B_{*}\in\mathbb{R}^{n\times m}$ and the process noise $\omega_{t+1}$ usually is assumed to be drawn i.i.d from a zero-mean gaussian distribution.
@@
With full knowledge of system parameters $A_{*}$, $B_{*}$, and assuming controllability of $(A_{*},B_{*})$, it is known that the optimal sequence of control inputs $u^{*}=\{u^{*}_{t}\}_{t=1}^{T}$ minimizing average expected cost
$J_{\mathcal{A}}=$ $\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}[\sum_{t=1}^{T}c_{t}]$ (3)
@@
In the absence of knowledge of the parameters $A_{*}$ and $B_{*}$, the core challenge is to design an algorithm that stabilizes the system while minimizing the accumulated quadratic cost, quantified by a finite-time performance measure. In this study, we adopt regret as the performance measure, defined as follows. For an arbitrary algorithm $\mathcal{A}$, which generates the input sequence $u=\{u_{t}\}_{t=1}^{T}$, regret is given by
$R_{\mathcal{A}}(T)=\sum_{t=1}^{T}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}-J_{*}\right),$ (4)
@@ 1.1 Contributions and Informal Statements of Results
The first algorithm proposed for the control of LQR with a worst-case regret bound of $\mathcal{O}(\sqrt{T})$ is the Optimism in the Face of Uncertainty (OFU)-based algorithm introduced by abbasi2011regret. This algorithm performs system identification by constructing confidence ellipsoids and designs the control policy by playing optimistically with respect to this set. Although the algorithm enjoys strong regret guarantees, its nonconvexity and the resulting computational burden has spurred subsequent research into more efficient formulations.
@@
Another property that may be important for making learning-based algorithms transferable across control domains is the provision of system-theoretic guarantees on the state-norm trajectory. Establishing such bounds, however, requires a thorough analysis of the closed-loop stability induced by the proposed algorithms. Generating individually stabilizing policies is not sufficient to ensure closed-loop stability, as instability (or state explosion) may still occur due to policy updates, a well known fact in literature on switched-systems control. Previous works have attempted to address this challenge through various approaches. For instance, certainty-equivalence (CE) methods simchowitz2020naive; jedra2022minimal employ a common Lyapunov function to guarantee Lyapunov stability of the closed-loop system. However, it is important to note that Lyapunov decay alone does not necessarily provide explicit high-probability, uniform control of the state norm in the presence of process noise. Under zero-mean Gaussian disturbances, Lyapunov decay guarantees only that the expected state norm, $\mathbb{E}\left[\|x_{t}\|^{2}\right]$, remains bounded, and deriving a high-probability bound requires additional effort. Although a high-probability state-norm bound is provided by simchowitz2020naive (see their Lemma 5.3 and its proof), this bound is conservative and does not accurately capture the true scale of possible state-norm excursions. Therefore, while Lyapunov stability is sufficient for controlling accumulated regret—since a finite number of such spikes can typically be absorbed in regret bounds—it is insufficient for obtaining high-probability guarantees on the state norm, which are essential for safety-critical systems. Another work that provides a high-probability state-norm bound is abeille2020efficient, but the bound is not fully system-theoretic and depends on a priori bounds, which may be loose. In another work, cohen2019learning introduced the notion of strong sequential stability, which enables an explicit high-probability guarantee on the state norm. This notion imposes a stricter requirement than standard stability: the algorithm must not only output stabilizing policies, but also ensure that any two consecutively generated policies remain close in a suitable sense, preventing the state norm from blowing up due to policy updates. While this approach provides a high-probability bound on the state norm, in this work we show that the strong sequential-stability requirement is conservative from the perspective of regret guarantees, leading to an $\mathcal{O}(\|P_{*}\|^{8})$ dependence on the operator norm of the DARE solution. This issue motivates us to propose a method that incorporates a mechanism inspired by dwell-time design in switched-systems control. Using this mechanism, we carefully regulate the exploration–exploitation trade-off to improve regret bounds while explicitly restricting state amplification and ensuring uniform control of the closed-loop trajectories. A natural first idea would be to require the algorithm to generate only stabilizing policies and then rely on a dwell-time condition to control both the state norm and the regret. However, sequential stability is not a binary property; rather, it forms a tunable continuum. Our analysis shows that fully enforcing sequential stability leads to suboptimal regret, while discarding it entirely and relying solely on a dwell-time–type design for exploration–exploitation tuning also degrades regret. The optimal performance arises at an intermediate relaxation level, where stability and exploration are balanced. We identify a specific tuning that guarantees an $\mathcal{O}(\|P_{*}\|^{6})$ dependence in the regret bound, while maintaining control of the state norm with a dependence of order $\mathcal{O}(\|P_{*}\|^{2.5})$. Our analysis reveals a clear trade-off between the state-norm bound and the regret bound in terms of their dependence on $\|P_{*}\|$, and it provides the flexibility to tune this trade-off as needed.
@@
Our core contributions are threefold: (i) We introduce new computationally efficient anytime-regret algorithms for LQ control, motivated by the need for methods that remain effective when the implementation horizon is not known a priori. (ii) We emphasize the importance of providing a high-probability, system-theoretic bound on the state norm. We offer a new theoretical understanding of strong sequential stability, highlight its limitations, and demonstrate how relaxing this notion can improve algorithm performance. A key component of this development is a novel mechanism, inspired by dwell-time design in switched-systems control, which enables principled tuning of the exploration–exploitation trade-off while ensuring both favorable regret bounds and closed-loop stability. (iii) We eliminate the need for a priori knowledge of the optimal average cost $J_{*}$, an assumption required by existing computationally efficient OFU-based algorithms.
@@
$(1)$ Any-time regret-guaranteed algorithm with system-theoretic bounds We propose an any-time, regret-guaranteed, SDP-based LQ control (ARSLO) algorithm built upon the SDP-based algorithm proposed by cohen2019learning, OSLO which is a horizon dependent algorithm. The OSLO algorithm when running for $T$ time steps requires $T$-dependent regularization parameters used for least square estimation and confidence bound construction as well as an initial parameter estimate in $\mathcal{O}(1/T^{1/4})$ neighborhood of true but unknown system dynamic parameters. The latter requires running a warm-up algorithm for $\mathcal{O}(\sqrt{T})$ time steps. To extend to a longer horizon using the doubling trick, such a warm-up phase must be implemented for the extended version in each epoch. Demonstrating this issue, we propose an any-time regret guaranteed algorithm by choosing an appropriate regularization parameter for the least square estimation and applying additive input perturbation $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, where $\|\Gamma_{t}\|_{op}=\mathcal{O}(1/\sqrt{t})$. Therefore, our ARSLO algorithm, provides anytime regret guarantees while also ensuring the strong sequential stability property of the generated policies (see Definition 1). Specifically, ARSLO achieves the regret bound
@@
Leaving aside the anytime regret guarantee of our proposed algorithms, their system-theoretic regret upper bound complements that of the Lag-LQR approach proposed by abeille2020efficient. Lag-LQR is developed by applying techniques from extended value iteration (EVI) to the Optimistic LQR with relaxed constraints, followed by a Lagrangian relaxation to design a computationally efficient algorithm. This algorithm comes with the following regret upper bound:
@@ 1.2 Related Works
In recent years, the machine learning community has shown increasing interest in addressing control design in the context of LQR, under various scenarios of what is known and what is unknown. Notably, agarwal2019online studied the setting with (nonstochastic) adversarial perturbations and established regret guarantees. cohen2018online investigated the control of linear systems with a known model but adversarially chosen quadratic costs, also providing regret guarantees. When the system model matrices are unknown, which is the main focus of this paper, existing efforts can be categorized into model-free and model-based settings. Model-free approaches design controllers without explicit system identification; instead, they minimize the cost function directly, relying solely on available data (see dean2020sample; fazel2018global; abbasi2019model). On the other hand, model-based approaches perform control design through explicit model identification or confidence set construction and typically provide regret guarantees. In this section, we briefly review the regret-guarantee literature that followed the OFU-based algorithm introduced by abbasi2011regret, which was the first algorithm for LQ control to achieve $\mathcal{O}(\sqrt{T})$ regret guarantees, albeit at a high computational cost.
@@
A commonly used categorization of regret-guaranteed algorithms for LQ control—also adopted in jedra2022minimal—classifies existing works into (i) self-tuning regulator–based approaches and (ii) OFU–based approaches. Self-tuning regulator algorithms estimate the unknown system matrices $(A_{*},B_{*})$, treat these estimates as the true parameters, solve the corresponding DARE, and design the control input accordingly. To ensure sufficient excitation of the system—and hence enable consistent estimation of $(A_{*},B_{*})$—the control inputs are typically perturbed with carefully tuned noise. OFU-based approaches, on the other hand, do not employ perturbation noise and instead design the control policy by acting optimistically with respect to a confidence set. Our algorithm, while OFU-based, also employs input perturbations for exploration. As a result, our approach can be viewed as lying at the intersection of these two categories. A more detailed discussion and quantitative comparisons are provided in Appendix A.
@@ 2 Problem Statement and Preliminaries
Notation. For matrix $M$, $\|M\|_{*}=\operatorname{tr}(\sqrt{M^{\top}M})$ is the trace norm, $\|M\|_{F}=\sqrt{\operatorname{tr}(M^{\top}M)}$ is the Frobenius norm and $\|M\|$ is the spectral (operator) norm. $\lambda_{\min}(.)$ represents the minimum eigenvalue. Finally, we use $A\bullet B$ to denote the entry-wise dot product between matrices, namely $A\bullet B=\operatorname{Tr}(A^{\top}B)$. We write $f\lesssim g$ to denote that $f(x)\leq C\,g(x)$ for some universal constant $C>0$. The notation $\mathcal{O}$ indicates dependencies only on the time $t$, the confidence parameter $\delta$, $\|P_{*}\|$, and the dimensions $n$ and $m$.
@@ 2.2 Preliminaries
Confidence Ellipsoid Construction To begin, we outline how a confidence ellipsoid around the true but unknown parameters $\Theta_{*}$ is constructed using the available measurements, leaving further details to the Appendix D. Let $X_{t}$ and $Z_{t}$ denote the matrices whose rows are $x_{2}^{\top},\ldots,x_{t}^{\top}$ and $z_{1}^{\top},\ldots,z_{t-1}^{\top}$, respectively. By applying least-square estimation, regularized with $\operatorname*{Tr}((\Theta-\Theta_{0})^{\top}(\Theta-\Theta_{0}))$ for some $\Theta_{0}$ and regularization parameter $\lambda$, the least-square estimate of $\Theta_{*}$ is given by:
@@
We denote its optimal solution by $P(\mathcal{C}_{t})$. Note that in cohen2018online, the term $\mu_{t}\|P\|_{*}V_{t}^{-1}$ on the right-hand side of their relaxed dual SDP introduces unnecessary looseness in the analysis and it can be replaced by $\mu_{t}\|P\|V_{t}^{-1}$. Our perturbation lemma (Lemma 7 in Appendix E), together with the corresponding lemma in their analysis, shows that such an SDP relaxation is not necessary.